V(s)50.94g/mol+O2g32.00g/mol----V2O3(s)149.88g/mol calculate the theoretical yield of vanadium (lll) oxide, assuming you begin with 200.00 grams vanadium metal
1
Expert's answer
2021-10-26T05:34:07-0400

4V + 3O2 → 2V2O3

m(V) "= \frac{200.0}{50.94} = 3.926 \;mol"

According to the reaction:

n(V2O3) "= \frac{1}{2}n(V) = \frac{1}{2} \times 3.926 = 1.963 \;mol"

m(V2O3) "= 1.963 \times 149.88 = 294.29 \;g"

Answer: 294.29 g